# MCGrad: Multicalibration at Web Scale

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Machine learning
@@ 2. Background on (Multi)Calibration
A predictor $f$ is perfectly calibrated if and only if $\mathbb{P}(Y=1\mid f(X)=p)=p$, where $p$ is the true underlying probability. Intuitively, for all input pairs $(x,y)\in D$, if $f$ predicts $0.8$, we expect that $80\%$ of them have 11 as label. Recently, Tygert (2023) introduced the Estimated Cumulative Calibration Error (ECCE), a novel parameter-free metric for measuring calibration that computes the maximum difference between the means of labels and predictions over any score interval:
@@
A common choice is to use a uniform bound on the MC-deviation, which is accommodated by choosing the scale parameter $\tau_{h}(f)$ to be a constant equal to 11. In Appendix A.1 we show that Definition 2.1, with $\tau_{h}(f)^{2}=\mathbb{E}\left[h(X)f(X)(1-f(X))\right]$, naturally matches with Multicalibration Error (MCE) by Guy et al. (2025), which quantifies multicalibration error by computing the maximum ECCE over the protected groups with a suitable normalization:
@@ 3.2. Fast Training and Prediction
GBDTs are regularized in multiple ways, including using a step size smaller than 11 for scaling each additional tree. While this is helps to avoid overfitting, it results in a predictor which can be strictly improved by rescaling it by a factor slightly greater than 11. As a result, the next round will attempt to apply this rescaling. Since linear rescaling is not easily expressed by decision trees, this may require many trees, ending up requiring unnecessarily many trees to achieve multicalibration.
@@
Note that this constant is typically very close to 11 and has a very limited detrimental effect on regularization. Rescaling is not the same as using a step size of 11: it affects the whole sum of trees (i.e., the whole predictor), while the step size only targets each tree sequentially (i.e., each weak learner).
$3.$ Preventing Overfitting
@@
Augmenting the data with the previous round’s model necessarily gives rise to regions of the augmented feature space that are particularly prone to overfitting. Consider a leaf of a partially constructed tree in GBDT. The leaf splitting algorithm can choose to split the leaf on values of the previous predictor $f_{t}(x)$. In this leaf, the left tail $f_{t}(x)<a$ for some small $a$ will contain only negative labels even if the true distribution may assign a positive probability to a positive label (and analogously for the right tail). In that case the new model can improve the likelihood simply by assigning a very low probability to this tail, as low as zero. Common regularization strategies, like setting the minimum number of samples per leaf, are insufficient to address this scenario. We solve this problem by using a more targeted form of regularization that limits the minimum total Hessian in a leaf, which is offered as one of multiple regularization techniques in LightGBM. The Hessian in a leaf $S$ equals $\sum_{(x,y)\in S}f_{t}(x)(1-f_{t}(x))$ and, as such, is a refined version of the simple sample size rule $|S|$. As the predicted probabilities become close to 0 or 11, the total Hessian becomes smaller and split in multiple leaves can no longer be considered. As shown in our ablation studies (Section 4) this regularization reduces overfitting and improves performance.
@@ 4.2. Experimental Results
Figure 1 (left-hand side) shows the Multicalibration Error (MCE) for all compared methods using all available features. Overall, MCGrad outperforms all baselines on $10$ out of the $11$ datasets by reducing the base predictor’s MCE between $4\%$ (Credit) and $93\%$ (ACSEmploy), and obtaining an average reduction of $56.1\%$. In contrast, DFMC, Isotonic, and HKRR only achieve an average reduction of, respectively, $11.9\%$, $9.6\%$, and $1.4\%$. In addition, MCGrad never harms the base predictor’s performance: even on MEPS, where there is low evidence of BasePred’s miscalibration (3.9 MCE), it slightly improves the MCE by $2.8\%$. On the other hand, Isotonic, DFMC, and HKRR increase (worsen) the MCE of BasePred on 33, 33 and 44 datasets respectively.
Table 1 reports the average performance over the datasets. MCGrad achieves by far the best average log loss improvement ($10.4\%$) and PRAUC improvement ($8.1\%$) over the base predictor. Besides these average improvements over datasets MCGrad never harms performance which helps build the necessary trust among ML engineers to delpoy it in production environments. Other methods don’t have this guarantee. For instance while, DFMC improves model performance metrics by a lower $1.8\%$ and $0.76\%$, it harms PRAUC and log loss on two datasets. Isotonic and HKRR even decrease the PRAUC on average by, respectively, $1\%$ and $5\%$ with a negligible change in log loss. For each metric, we rank all compared methods from 11 (best) to 55 (worst) and report the average ranking position in Table 1. A lower average rank indicates how often the method is preferred, as opposed to the average performance, which reflects the magnitude. MCGrad consistently ranks as #1\#1 baseline on log loss and PRAUC, while only ranking third for MCE on the MEPS dataset. Finally, MCGrad ranks second best method in terms of global calibration (ECCE), losing against Isotonic Regression, which is expected because it is trained to directly optimize the global calibration of the base predictor.
@@
To verify that good performance on the MCE metric for all groups implies also protection for prespecified groups, we report results just for these groups in Figure 1 (right-hand side). Overall, MCGrad outperforms all baselines on 55 out of $11$ datasets achieving a $54\%$ average MCE improvement, and is followed by DFMC that is the best baseline on three datasets (ACSMobility, HDMA, ACSTravel) and achieves an average MCE improvement of $51\%$. HKRR obtains an average improvement in MCE of $36\%$, while Isotonic only achieves an improvement of $16\%$. Note that HKRR and DFMC have direct access to these groups, while MCGrad only has access to the features used to define the groups.
@@
Because MCGrad recurrently improves its own predictions by using $T$ rounds, one relevant question is whether $T>1$ is really necessary in practice. For this goal, we compare MCGrad with its variant Only One Round that forces the number of rounds T=1T=1.333Early stopping can possibly return $T=0$. Figure 2 (purple line) shows the relative improvement of Only One Round for the four considered metrics and all benchmark datasets. In the majority of cases, it is evident that Only One Round deteriorates MCGrad’s performance: on the MCE and ECCE axes, it more than doubles its value (relative improvement <100<-100) for, respectively, 66 and 55 datasets, while, on the log loss and PRAUC axes, it worsens the performance by approximately $2\%$ or more on, respectively, 88 and 66 out of $11$ datasets. Interestingly, the largest deterioration of performance corresponds to MCGrad running for the highest number of rounds: the optimal $T$ for ACSEmploy, ACSIncome, ACSPublic, and ACSMobility is, respectively, 2020, $35$, $27$ and $22$.
@@
Rescaling the logits of all predictions by learning a constant speeds up MCGrad training by reducing rounds, especially on very large datasets. As a downside, it might affect MCGrad’s performance. Figure 2 (light blue bar) shows the relative deterioration after removing rescaling. Effects are mild and balanced: removal worsens MCE, log loss, and PRAUC on 66, 66, and 77 datasets respectively, with average drops of $9\%$, $0.5\%$, and $0.03\%$.
LightGBM’s default min sum Hessian leaf (MSHL) is $0.001$ with a min leaf sample size of 2020. We claim that higher MHL values prevent probabilities from pushing to extremes, reducing overfitting (Section 3.3). We compare MCGrad with MHL=0.001=0.001 versus our default of 2020. Figure 2 (dark blue bar) shows mild deterioration using $0.001$, likely due to limited dataset sizes. On average, lower MHL slightly harms performance, but this becomes evident on ACSEmploy (the largest dataset), where the MHL variant runs only 22 rounds (vs. $10$) because our regularization mechanism stops the training earlier to avoid overfitting.
@@ A.2.2. A Bound on the Multicalibration Error
The proposition provides a bound on the miscalibration error in terms of the gap between successive predictor values $F_{T+1}$ and $F_{T}$, $\epsilon_{T}$ that captures the gap between the loss function values achieved by the ensemble of regression trees added by the GBM and the optimal ensemble in round $T$, and $C_{T}$ that depends on how much the predictor $f_{T}$ concentrates around the boundary values 0 and 11. The dependence on the gap between $F_{T}$ and $F_{T+1}$ is such that the multicalibration error decreases as the gap between $F_{T}$ and $F_{T+1}$ decreases. The error term $\epsilon_{T}$ can be controlled by the number of iterations $M_{t}$ used by GBM in round $t$ of the algorithm, as demonstrated in Equation (5).

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 2. Background on (Multi)Calibration
A predictor $f$ is perfectly calibrated if and only if $\mathbb{P}(Y=1\mid f(X)=p)=p$, where $p$ is the true underlying probability. Intuitively, for all input pairs $(x,y)\in D$, if $f$ predicts $0.8$, we expect that $80\%$ of them have $1$ as label. Recently, Tygert (2023) introduced the Estimated Cumulative Calibration Error (ECCE), a novel parameter-free metric for measuring calibration that computes the maximum difference between the means of labels and predictions over any score interval:
@@
A common choice is to use a uniform bound on the MC-deviation, which is accommodated by choosing the scale parameter $\tau_{h}(f)$ to be a constant equal to $1$. In Appendix A.1 we show that Definition 2.1, with $\tau_{h}(f)^{2}=\mathbb{E}\left[h(X)f(X)(1-f(X))\right]$, naturally matches with Multicalibration Error (MCE) by Guy et al. (2025), which quantifies multicalibration error by computing the maximum ECCE over the protected groups with a suitable normalization:
@@ 3.2. Fast Training and Prediction
GBDTs are regularized in multiple ways, including using a step size smaller than $1$ for scaling each additional tree. While this is helps to avoid overfitting, it results in a predictor which can be strictly improved by rescaling it by a factor slightly greater than $1$. As a result, the next round will attempt to apply this rescaling. Since linear rescaling is not easily expressed by decision trees, this may require many trees, ending up requiring unnecessarily many trees to achieve multicalibration.
@@
Note that this constant is typically very close to $1$ and has a very limited detrimental effect on regularization. Rescaling is not the same as using a step size of $1$: it affects the whole sum of trees (i.e., the whole predictor), while the step size only targets each tree sequentially (i.e., each weak learner).
3.3. Preventing Overfitting
@@
Augmenting the data with the previous round’s model necessarily gives rise to regions of the augmented feature space that are particularly prone to overfitting. Consider a leaf of a partially constructed tree in GBDT. The leaf splitting algorithm can choose to split the leaf on values of the previous predictor $f_{t}(x)$. In this leaf, the left tail $f_{t}(x)<a$ for some small $a$ will contain only negative labels even if the true distribution may assign a positive probability to a positive label (and analogously for the right tail). In that case the new model can improve the likelihood simply by assigning a very low probability to this tail, as low as zero. Common regularization strategies, like setting the minimum number of samples per leaf, are insufficient to address this scenario. We solve this problem by using a more targeted form of regularization that limits the minimum total Hessian in a leaf, which is offered as one of multiple regularization techniques in LightGBM. The Hessian in a leaf $S$ equals $\sum_{(x,y)\in S}f_{t}(x)(1-f_{t}(x))$ and, as such, is a refined version of the simple sample size rule $|S|$. As the predicted probabilities become close to 0 or $1$, the total Hessian becomes smaller and split in multiple leaves can no longer be considered. As shown in our ablation studies (Section 4) this regularization reduces overfitting and improves performance.
@@ 4.2. Experimental Results
Figure 1 (left-hand side) shows the Multicalibration Error (MCE) for all compared methods using all available features. Overall, MCGrad outperforms all baselines on $10$ out of the $11$ datasets by reducing the base predictor’s MCE between $4\%$ (Credit) and $93\%$ (ACSEmploy), and obtaining an average reduction of $56.1\%$. In contrast, DFMC, Isotonic, and HKRR only achieve an average reduction of, respectively, $11.9\%$, $9.6\%$, and $1.4\%$. In addition, MCGrad never harms the base predictor’s performance: even on MEPS, where there is low evidence of BasePred’s miscalibration (3.9 MCE), it slightly improves the MCE by $2.8\%$. On the other hand, Isotonic, DFMC, and HKRR increase (worsen) the MCE of BasePred on $3$, 33 and $4$ datasets respectively.
Table 1 reports the average performance over the datasets. MCGrad achieves by far the best average log loss improvement ($10.4\%$) and PRAUC improvement ($8.1\%$) over the base predictor. Besides these average improvements over datasets MCGrad never harms performance which helps build the necessary trust among ML engineers to delpoy it in production environments. Other methods don’t have this guarantee. For instance while, DFMC improves model performance metrics by a lower $1.8\%$ and $0.76\%$, it harms PRAUC and log loss on two datasets. Isotonic and HKRR even decrease the PRAUC on average by, respectively, $1\%$ and $5\%$ with a negligible change in log loss. For each metric, we rank all compared methods from $1$ (best) to $5$ (worst) and report the average ranking position in Table 1. A lower average rank indicates how often the method is preferred, as opposed to the average performance, which reflects the magnitude. MCGrad consistently ranks as #1\#1 baseline on log loss and PRAUC, while only ranking third for MCE on the MEPS dataset. Finally, MCGrad ranks second best method in terms of global calibration (ECCE), losing against Isotonic Regression, which is expected because it is trained to directly optimize the global calibration of the base predictor.
@@
To verify that good performance on the MCE metric for all groups implies also protection for prespecified groups, we report results just for these groups in Figure 1 (right-hand side). Overall, MCGrad outperforms all baselines on $5$ out of $11$ datasets achieving a $54\%$ average MCE improvement, and is followed by DFMC that is the best baseline on three datasets (ACSMobility, HDMA, ACSTravel) and achieves an average MCE improvement of $51\%$. HKRR obtains an average improvement in MCE of $36\%$, while Isotonic only achieves an improvement of $16\%$. Note that HKRR and DFMC have direct access to these groups, while MCGrad only has access to the features used to define the groups.
@@
Because MCGrad recurrently improves its own predictions by using $T$ rounds, one relevant question is whether $T>1$ is really necessary in practice. For this goal, we compare MCGrad with its variant Only One Round that forces the number of rounds T=1T=1.333Early stopping can possibly return $T=0$. Figure 2 (purple line) shows the relative improvement of Only One Round for the four considered metrics and all benchmark datasets. In the majority of cases, it is evident that Only One Round deteriorates MCGrad’s performance: on the MCE and ECCE axes, it more than doubles its value (relative improvement <100<-100) for, respectively, $6$ and $5$ datasets, while, on the log loss and PRAUC axes, it worsens the performance by approximately $2\%$ or more on, respectively, $8$ and $6$ out of $11$ datasets. Interestingly, the largest deterioration of performance corresponds to MCGrad running for the highest number of rounds: the optimal $T$ for ACSEmploy, ACSIncome, ACSPublic, and ACSMobility is, respectively, $20$, $35$, $27$ and $22$.
@@
Rescaling the logits of all predictions by learning a constant speeds up MCGrad training by reducing rounds, especially on very large datasets. As a downside, it might affect MCGrad’s performance. Figure 2 (light blue bar) shows the relative deterioration after removing rescaling. Effects are mild and balanced: removal worsens MCE, log loss, and PRAUC on $6$, 66, and $7$ datasets respectively, with average drops of $9\%$, $0.5\%$, and $0.03\%$.
LightGBM’s default min sum Hessian leaf (MSHL) is $0.001$ with a min leaf sample size of $20$. We claim that higher MHL values prevent probabilities from pushing to extremes, reducing overfitting (Section 3.3). We compare MCGrad with MHL=0.001=0.001 versus our default of $20$. Figure 2 (dark blue bar) shows mild deterioration using $0.001$, likely due to limited dataset sizes. On average, lower MHL slightly harms performance, but this becomes evident on ACSEmploy (the largest dataset), where the MHL variant runs only $2$ rounds (vs. $10$) because our regularization mechanism stops the training earlier to avoid overfitting.
@@ A.2.2. A Bound on the Multicalibration Error
The proposition provides a bound on the miscalibration error in terms of the gap between successive predictor values $F_{T+1}$ and $F_{T}$, $\epsilon_{T}$ that captures the gap between the loss function values achieved by the ensemble of regression trees added by the GBM and the optimal ensemble in round $T$, and $C_{T}$ that depends on how much the predictor $f_{T}$ concentrates around the boundary values 0 and $1$. The dependence on the gap between $F_{T}$ and $F_{T+1}$ is such that the multicalibration error decreases as the gap between $F_{T}$ and $F_{T+1}$ decreases. The error term $\epsilon_{T}$ can be controlled by the number of iterations $M_{t}$ used by GBM in round $t$ of the algorithm, as demonstrated in Equation (5).